OpenStudy (anonymous):

Find a cubic function with the given zeros. 6, -5, 2

OpenStudy (anonymous):

If a , b and c are the three zeros of a cubic polynomial with the variable x then you can find out the equation by (x-a) ( x-b) ( x - c ) = 0

OpenStudy (anonymous):

give it a try @melacho ... try to solve it now

OpenStudy (kirbykirby):

ok you beat me to it lol

OpenStudy (anonymous):

i still dont get it :>

OpenStudy (anonymous):

haha @kirbykirby

OpenStudy (anonymous):

@melacho ... the given zeros are 6 , -5 , 2 and let 6 be a... i.e. 6 = a -5 be b ...i.e. -5 = b and 2 be c ..... i.e. 2 = c Now... try to use these values in (x-a) ( x-b) ( x - c ) = 0

OpenStudy (anonymous):

(x-6) (x-(-5)) (x-2)=0?

OpenStudy (anonymous):

yup..!

OpenStudy (agent0smith):

You want a function, so just make it f(x) = (x-6) (x+5) (x-2)

OpenStudy (anonymous):

but these are my choices f(x) = x3 - 3x2 + 28x + 60 f(x) = x3 - 3x2 - 28x - 60 f(x) = x3 + 3x2 - 28 + 60 f(x) = x3 - 3x2 - 28x + 60

OpenStudy (agent0smith):

Multiply it out.

OpenStudy (kirbykirby):

If you're not sure where this idea is coming from, you can try googling "Factor Theorem". It essentially states that if "a" is a zero of a polynomial, then (x-a) is a factor of the polynomial. Hence, that is why @Shizen suggested using (x-a)(x-b)(x-c)

OpenStudy (anonymous):

Multipy any of the two factors first and then multiply the result with the third one....

OpenStudy (agent0smith):

f(x) = (x-6) (x+5) (x-2) first multiply (x-6) (x+5) then multiply that result by (x-2)

OpenStudy (anonymous):

f(x) = x3 - 3x2 - 28x - 60

OpenStudy (anonymous):

well.. just recalulate the constant term....

OpenStudy (anonymous):

f(x) = x3 - 3x2 + 28x + 60

OpenStudy (anonymous):

Can you show me, the way you are solving it ?? cos you are commiting a mistake....

OpenStudy (anonymous):

f(x) = x3 + 3x2 - 28 + 60

OpenStudy (anonymous):

i did:

OpenStudy (anonymous):

x3 + 3x2 - 28 + 60 well you missed the x after -28...:D...

OpenStudy (anonymous):

f(x) = x3 - 3x2 - 28x + 60

OpenStudy (anonymous):

ahaan,..! thats correct....:D